A university bookstore recently sold a wirebound graph-paper notebook for $2.15, and a college-ruled notebook for $3.43. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $134.38. How many of each type were sold?
Answer by mananth(16946)

A university bookstore recently sold a wirebound graph-paper notebook for $2.15, and a college-ruled notebook for $3.43. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $134.38. How many of each type were sold?
...
x graph books
y ruled books
..
x+y = 50................1
2.15x+3.43y = 134.38.....2
...
multiply (1) by -2.15
-2.15x-2.15y=-50*2.15
add this to (2)
3.43y-2.15y = 134.38-107.50
1.28 y= 26.88
y = 21 ruled books
29 graph books
...
m.ananth@hotmail.ca